# Visual Program Distillation: Distilling Tools and Programmatic Reasoning into Vision-Language Models

### 05 Dec 2023 v1

@@ Abstract
Figure 1: We introduce Visual Program Distillation (VPD), a training framework which leverages LLM-generated programs to synthesis multimodal chain-of-thought training data for Vision-Language Models (VLMs). Our generalist models trained with VPD, PaLI-3-VPD (5B) and PaLI-X-VPD (55B), outperform prior VLMs on a broad range of tasks, while producing human-interpretable and faithful reasoning steps.
@@ 1 Introduction
Vision-language models (VLMs) have become the pretrained backbone for many computer vision tasks [2, 59, 34, 39, 78, 4, 72, 62, 7, 41, 9, 10, 11]. Yet, all these models still fall short of solving numerous visual reasoning tasks expected of competent vision models. Even state-of-the-art (SOTA) proprietary vision-language models such as GPT-4V [49] do not do well at tasks that involve counting and spatial reasoning [69]. They find it difficult to count (TallyQA [1]), to compositionally reason (GQA [26]), and to reason with external knowledge (OK-VQA [44], A-OKVQA [53]). Many of these tasks require VLMs to conduct compositional reasoning, which still remains an unsolved challenge. For instance, answering the question “Who invented the musical instrument on the right?” involves a composition of skills: identifying objects, applying spatial reasoning to locate the one on the right, recognizing the musical instrument, and accessing prior knowledge to retrieve the inventor.
@@
Another line of work is visual instruction tuning [39], which tries to distill the instruction-following ability of LLMs into VLMs. They prompt LLMs with image captions and bounding box annotations in order to generate queries and answers that are used to fine-tune VLMs [39, 64, 7, 23]. However, this approach has important limitations: image captions can miss fine-grained visual information, and LLM are prone to produce inconsistent outputs for custom vision representations like bounding boxes [18]. As a result, existing instruction-tuned VLMs still struggle with tasks that requires complex visual reasoning [40, 38, 14, 7].
In this work, we present Visual Program Distillation (VPD), a novel distillation method that induces LLM-like complex reasoning capabilities into vision-language models (Fig. 2). As the name suggests, VPD combines two key insights to deliver a training paradigm that surpasses the sum of its parts: It relies on (1) advancements in visual programs that use tools [20] and (2) the recent breakthroughs in distillation through chain-of-thought reasoning [21]. Given a labeled training dataset of complex visual tasks, VPD generates a correct program, and then distills its reasoning steps into vision-language models. To avoid using programs that give the wrong answer, VPD prompts an LLM to generate multiple candidate programs, and executes every one of them. When labeled data is available, it then filters for programs that produce the correct answer upon execution. Therefore, our programs comprise multiple vision tools, are executable, and yield the correct answer when executed. Next, VPD rewrites the correct programs as natural language chain-of-thought instructions and uses step-by-step distillation [21] to inject the reasoning abilities into VLMs.
@@ 2 Related work
Most recent generative VLMs share a common structure: a pre-trained visual encoder, a pre-trained LLM, and a connector between the two modalities [2, 9, 59, 34, 39, 78, 4, 72, 62, 41]. The models are trained on large-scale image-text pairs from various tasks to adapt to both modalities. They are also tuned on LLM-generated visual instructions to enable models to follow versatile instructions from users [39]. Some models also include bounding boxes in the pre-training data to improve the VLM on visually grounded tasks [77, 50, 60, 7, 41, 6, 10]. The bounding boxes are usually retrieved from in COCO [37] and Visual Genome [32]. Different from prior work, our method does not rely on provided dense annotations. We use LLM-generated code and specialized vision tools to generate our own visual instruction data.
@@ 3.1 Program generation and verification
    Converting program execution traces into chains-of-thought: We rewrite $t$ into a CoT $c$ using an LLM.
@@ 5 Experiments on content moderation
Prior experiments focus on training generalist VLMs. Here, we explore the effectiveness of VPD at quickly adapting models to real-world applications from a different domain than the training data. We experiment on Hateful Memes [29], a content moderation dataset where the task is to classify if a meme contains hateful content. The target labels are “yes” or “no”, and models are evaluated in terms of classification accuracy and AUC-ROC. We experiment with two settings: supervised, in which the models are trained on the provided training set with $8,500$ labels, and unsupervised, in which no human labels are provided. We provide qualitative examples of the rationales our models give in Figure 7.
@@
Table 3: Results on Hateful Memes [29] seen test set. We improve SOTA for both supervised and unsupervised settings. Surprisingly, unsupervised PaLI-X-VPD outperforms supervised VisualBERT.
@@ 6 Limitations and Future Directions
We find that programs struggles in complex scenarios when objects are occluded, or the scene contains too many objects. Adding dense-labeling tools like Segment Anything [31] may address this challenge. Other recent work (e.g., LISA [33]) have proposed combining segmentation with LLMs. Future work can make dense labeling tools available in VPD in a similar way, which will further boost VLM performance.
@@ 8 Acknowledgements
We thank Tanmaya Dabral, Jialin Wu, Radu Soricut, Aditya Avinash, Howard Zhou, Jiao Sun, Deqing Fu, Cyrus Rashtchian, Andrew Tomkins, and Tom Duerig for helpful project discussions and technical support in this project.
Appendix A Examples of our Data-Synthesis Pipeline
We first present step-by-step examples of our data-synthesis pipeline. As discussed in §3, the pipeline contains four steps: program generation, program execution, program filtering, and converting execution traces into chains-of-thought. Here we show three filtered examples from three datasets we use. As shown, LLMs like PaLM-2 [3] is able to generate complex programs to solve compositional vision tasks. The synthesized CoTs are used to fine-tune our VLMs.
@@ Appendix B Experimental Details
B.1 Details of Model Architectures
For both PaLI models, the architectures are similar: images are encoded into visual tokens individually via a visual encoder. Then, the vision tokens along with the textual inputs are passed through an encoder-decoder UL2 Transformer [57]. The PaLI models were then pre-trained with image-text pairs to perform multimodal tasks. Specifically, PaLI-3 [11] uses a pre-trained 2B SigLIP [76] as visual encoder, and a 3B UL2. The image resolution is $812\times 812$. PaLI-X [10] uses a pre-trained VIT-22B [15] as visual encoder, and a 32B UL2. The image resolution is $756\times 756$. Please refer to the PaLI-3 [11] and PaLI-X [10] papers about more architecture details.
B.2 Details of Data
@@ B.3 Training Details
We use LoRA [22] to fine-tune both PaLI-3 [11] and PaLI-X [10]. For generalist training, we add LoRA weights on each linear layer in the attention blocks and multilayer perceptron (MLP) blocks for both the encoder and decoder in the UL2 transformer. For both models, we use $rank=8$. We use a cosine learning rate schedule, with warm-up ratio 1% and peak learning rate $41e-$ 4. For all models and all settings, we use a batch size of 128 and fine-tune the pre-trained model for 8,000 steps. In terms of training time, we train PaLI-X-VPD with 128 TPU-v3 [27] and it takes about 2 days to finish training. For PaLI-3-VPD, we use 32 TPU-v4 and training takes about 20 hours. We still observe a steady loss drop when we terminate training, which indicates that more computation may lead to even better performance. For per-task fine-tuning, to avoid overfitting, we reduce the number of training parameters. For both models, we only add LoRA weights to encoder layers. We use LoRA $rank=4$ for PaLI-X-VPD and $rank=8$ for PaLI-3-VPD. The peak learning rate is $41e-$ 4 and we use cosine learning schedule, with warmup ratio 1%. For all per-task fine-tuning experiments, we use a batch size of 64. We train for 1 epoch on GQA, and 3 epochs on all other datasets. We use the AdamW [30] optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.98$, and bfloat16 for all experiments.
@@ Appendix C Human Evaluation
We asked our human annotators to first evaluate each model’s answer, using the criteria described in §4.3. After rating each model answer separately, we also asked them to choose a preferred answer among the two. However, we observed that there are cases where one or both models have similar answers, or both answers are incorrect so it would be difficult to choose a favorite, so we allowed annotators to also choose “Both” or “Neither”, giving the annotators the following instruction: “Please try to choose "Answer 1 is better" or "Answer 2 is better" whenever possible. We also give you the option to choose "Both are equally good." or "Both are too bad to make a choice." for the cases when it is hard to make a choice either because both answers are correct and similar, or because both answers are wrong so it makes no sense to choose a favorite.”
We show some examples from our human evaluation in Table 6. The table contains the images and corresponding text queries (column 2), the answers provided by the two models we compared—PaLI-X Instruct (column 3) and PaLI-X-VPD (column 4)—along with the corresponding annotations given by the human annotators. The human annotations are aggregated across 3 raters per sample. Finally, column 5 shows which of the two answers was preferred by the human raters. When a model’s answer includes a bounding box, we annotate it on the image for convenience.

### 05 Apr 2024 v2

@@ Abstract
Figure 1: We introduce Visual Program Distillation (VPD), a framework which leverages tools and LLM-generated programs to synthesize cross-modal reasoning data for vision-language model (VLM) training. VPD distills the skills in specialist models (e.g., detection, depth estimation) and LLM reasoning ability into VLMs. Our generalist models trained with VPD, PaLI-X-VPD (55B), outperform prior VLMs on a broad range of tasks, while producing human-interpretable and faithful reasoning steps.
@@ 1 Introduction
Vision-language models (VLMs) have become the pretrained backbone for many computer vision tasks [2, 60, 34, 39, 79, 4, 73, 63, 7, 41, 9, 10, 11]. Yet, all these models still fall short of solving numerous visual reasoning tasks expected of competent vision models. Even state-of-the-art (SOTA) proprietary vision-language models such as GPT-4V [49] do not perform well on tasks that involve counting and spatial reasoning [70]. They find it difficult to count (TallyQA [1]), to compositionally reason (GQA [26]), and to reason with external knowledge (OK-VQA [44], A-OKVQA [53]). Many of these tasks require VLMs to conduct compositional reasoning, which still remains an unsolved challenge. For instance, answering the question “Who invented the musical instrument on the right?” involves a composition of skills: identifying objects, applying spatial reasoning to locate the one on the right, recognizing the musical instrument, and accessing prior knowledge to retrieve the inventor.
@@
Another line of work is visual instruction tuning [39], which tries to distill the instruction-following ability of LLMs into VLMs. They prompt LLMs with image captions and bounding box annotations in order to generate queries and answers that are used to fine-tune VLMs [39, 65, 7, 23]. However, this approach has important limitations: image captions can miss fine-grained visual information, and LLM are prone to producing inconsistent outputs for custom vision representations like bounding boxes [18]. As a result, existing instruction-tuned VLMs still struggle with tasks that requires complex visual reasoning [40, 38, 14, 7].
In this work, we present Visual Program Distillation (VPD), a novel distillation method that induces complex cross-modal reasoning capabilities into vision-language models (Fig. 2). As the name suggests, VPD combines two key insights to deliver a training paradigm that surpasses the sum of its parts: It relies on (1) advancements in visual programs that use tools [20] and (2) the recent breakthroughs in distillation through chain-of-thought reasoning [21]. Given a labeled training dataset of complex visual tasks, VPD generates a correct program, and then distills its reasoning steps into vision-language models. To avoid using programs that give the wrong answer, VPD prompts an LLM to generate multiple candidate programs, and executes every one of them. When labeled data is available, it then filters for programs that produce the correct answer upon execution. Therefore, our programs comprise multiple vision tools, are executable, and yield the correct answer when executed. Next, VPD rewrites the correct programs as natural language chain-of-thought instructions and uses step-by-step distillation [21] to inject the reasoning abilities into VLMs.
@@ 2 Related work
Most recent generative VLMs share a common structure: a pre-trained visual encoder, a pre-trained LLM, and a connector between the two modalities [2, 9, 60, 34, 39, 79, 4, 73, 63, 41]. The models are trained on large-scale image-text pairs from various tasks to adapt to both modalities. They are also tuned on LLM-generated visual instructions to enable models to follow versatile instructions from users [39]. Some models also include bounding boxes in the pre-training data to improve the VLM on visually grounded tasks [78, 50, 61, 7, 41, 6, 10]. The bounding boxes are usually retrieved from COCO [37] and Visual Genome [32]. Different from prior work, our method does not rely on provided dense annotations. We use LLM-generated code and specialized vision tools to generate our own visual instruction-tuning data.
@@ 3.1 Program generation and verification
    Converting program execution traces into chain-of-thought rationales: We rewrite $t$ into a CoT $c$ using an LLM.
@@ 5 Experiments on content moderation
Prior experiments focus on training generalist VLMs. Here, we explore the effectiveness of VPD at quickly adapting models to real-world applications from a different domain than the training data [e.g., 56]. We experiment on Hateful Memes [29], a content moderation dataset where the task is to classify if a meme contains hateful content. The target labels are “yes” or “no”, and models are evaluated in terms of classification accuracy and AUC-ROC. We experiment with two settings: supervised, in which the models are trained on the provided training set with $8,500$ labels, and unsupervised, in which no human labels are provided. We provide qualitative examples of the rationales our models give in Figure 7.
@@
Table 3: Results on Hateful Memes [29] seen test set. We improve SOTA for both zero-shot and supervised settings. Unsupervised PaLI-X-VPD outperforms strong supervised baselines.
@@ 6 Limitations and Future Directions
We find that programs struggles in complex scenarios when objects are occluded, or the scene contains too many objects. Adding dense-labeling tools like Segment Anything [31] may address this challenge. For example, recently LISA [33] have proposed combining segmentation with LLMs. Future work can make dense labeling tools available in VPD in a similar way, which will further boost VLM performance.
@@ 8 Acknowledgements
We thank Tanmaya Dabral, Jialin Wu, Radu Soricut, Aditya Avinash, Howard Zhou, Jiao Sun, Deqing Fu, Cyrus Rashtchian, Andrew Tomkins, Tom Duerig, as well as the Agile Modeling team for helpful project discussions and technical support in this project.
Appendix A Examples outputs of our data-synthesis pipeline
We first present step-by-step examples of our data-synthesis pipeline. As discussed in §3, the pipeline contains four steps: program generation, program execution, program filtering, and converting execution traces into chains-of-thoughts. Here we show three filtered examples from three datasets we use. As shown, LLMs like PaLM-2 [3] are able to generate complex programs to solve compositional vision tasks. The synthesized CoTs are used to fine-tune our VLMs.
@@ Appendix B Experimental details
B.1 Model architecture details
For both PaLI models that we considered in our experiments, the architectures are similar: images are encoded into visual tokens individually via a visual encoder. Then, the vision tokens along with the textual inputs are passed through an encoder-decoder UL2 Transformer [58]. The PaLI models were then pre-trained with image-text pairs to perform multimodal tasks. Specifically, PaLI-3 [11] uses a pre-trained 2B SigLIP [77] as visual encoder, and a 3B UL2. The image resolution is $812\times 812$. PaLI-X [10] uses a pre-trained VIT-22B [15] as visual encoder, and a 32B UL2. The image resolution is $756\times 756$. Please refer to the PaLI-3 [11] and PaLI-X [10] papers about more architecture details.
B.2 Datasets
@@ B.3 Training details
We use LoRA [22] to fine-tune both PaLI-3 [11] and PaLI-X [10]. For generalist training, we add LoRA weights on each linear layer in the attention blocks and multilayer perceptron (MLP) blocks for both the encoder and decoder in the UL2 transformer. For both models, we use $rank=8$. We use a cosine learning rate schedule, with warm-up ratio $1\%$ and peak learning rate $41e-$ 4. For all models and all settings, we use a batch size of $128$ and fine-tune the pre-trained model for $8,000$ steps. In terms of training time, we train PaLI-X-VPD with 128 TPU-v3 [27] and it takes about $2$ days to finish training. For PaLI-3-VPD, we use $32$ TPU-v4 and training takes about $20$ hours. We still observe a steady loss drop when we terminate training, which indicates that more computation may lead to even better performance. For per-task fine-tuning, to avoid overfitting, we reduce the number of training parameters. For both models, we only add LoRA weights to encoder layers. We use LoRA $rank=4$ for PaLI-X-VPD and $rank=8$ for PaLI-3-VPD. The peak learning rate is $41e-$ 4 and we use a cosine learning schedule, with warm-up ratio $1\%$. For all per-task fine-tuning experiments, we use a batch size of $64$. We train for $1$ epoch on GQA, and $3$ epochs on all other datasets. We use the AdamW [30] optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.98$, and bfloat16 for all experiments.
B.4 Inference costs
We sampled $300$ questions and measured the computation cost. Using 128 TPU v5, code generation takes on average $4.7s$, while program execution takes $4.2s$. With the same resources, PaLI-X-VPD takes $0.8s$ per question. The cost gap ($0.8s$ vs $8.9s$) is very large with immense consequences for practical applications.
@@ Appendix C Human evaluation
We asked our human annotators to first evaluate each model’s answer, using the criteria described in §4.3. After rating each model answer separately, we also asked them to choose a preferred answer between the two. However, we observed that there are cases where one or both models have similar answers, or both answers are incorrect, in which case it would be difficult for the annotators to choose a favorite, so we also provided the annotators with the options “Both” or “Neither”, giving them the following instruction: “Please try to choose "Answer 1 is better" or "Answer 2 is better" whenever possible. We also give you the option to choose "Both are equally good." or "Both are too bad to make a choice." for the cases when it is hard to make a choice either because both answers are correct and similar, or because both answers are wrong so it makes no sense to choose a favorite.”
We show some examples from our human evaluation in Table 6. The table contains the images and corresponding text queries (column 2), the answers provided by the two models we compared—PaLI-X Instruct (column 3) and PaLI-X-VPD (column 4)—along with the corresponding annotations given by the human annotators. The human annotations are aggregated across 3 raters per sample. Finally, column 5 shows which of the two answers was preferred by the human raters. When a model’s answer includes a bounding box, we annotate it on the image for convenience. Examples are as follow: